# Typicalness-Aware Learning for Failure Detection

### v1

@@ 4.1 Comparisions with the State-of-the-art on CIFAR
ResNet110 He et al. (2016a) MSPHendrycks and Gimpel (2016) 99.83 67.49 84.07 293.44 83.41 74.55 376.42 66.92 84.00 72.01
Cosine Zhang and Xiang (2023) 96.53 65.15 84.42 271.13 78.30 79.31 361.87 56.23 86.93 72.01
Energy Liu et al. (2020) 135.85 74.66 77.20 275.39 83.18 77.78 387.44 66.96 83.21 72.01
MaxLogitHendrycks and Gimpel (2016) 133.19 72.33 77.96 275.85 82.53 77.73 385.81 65.08 83.56 72.01
Entropy Tian et al. (2022a) 100.05 66.28 84.12 287.62 81.20 75.93 373.49 61.33 84.73 72.01
Mahalanobis De Maesschalck et al. (2000) 114.21 73.48 80.41 263.49 72.70 80.55 368.55 58.74 85.74 72.01
Gradnorm Huang et al. (2021) 369.86 98.82 35.30 490.21 98.17 49.26 679.48 98.69 42.76 72.01
SIRC Xia and Bouganis (2022) 100.56 66.37 84.01 287.93 81.03 75.90 374.12 61.29 84.65 72.01
LogitNorm Wei et al. (2022) 125.59 72.87 79.71 235.50 73.23 83.35 356.88 55.80 87.80 70.34
OpenMix Zhu et al. (2023a) 85.66 63.82 85.25 342.16 87.03 69.27 406.80 70.37 80.25 73.68
TAL 98.40 66.57 83.55 260.84 78.84 80.07 353.66 55.90 87.08 72.45
FMFP Zhu et al. (2022) 69.83 62.17 87.15 284.13 81.77 74.98 345.37 62.99 84.86 75.18
TAL w/ FMFP 74.74 67.99 84.82 224.72 64.66 86.00 310.57 46.33 90.48 75.59
WRNet Zagoruyko and Komodakis (2016) MSP Hendrycks and Gimpel (2016) 57.60 60.13 87.47 264.67 80.04 79.41 321.39 57.26 87.23 78.55
Cosine Zhang and Xiang (2023) 56.68 58.65 87.56 279.53 81.60 77.91 334.04 57.96 86.08 78.55
Energy Liu et al. (2020) 66.33 65.64 84.73 257.78 78.93 80.83 323.27 58.70 87.07 78.55
MaxLogitHendrycks and Gimpel (2016) 64.85 62.70 85.36 258.57 79.35 80.67 322.42 56.89 87.29 78.55
Entropy Tian et al. (2022a) 59.31 62.78 86.80 260.92 79.50 80.26 320.66 56.92 87.40 78.55
Mahalanobis De Maesschalck et al. (2000) 76.43 72.99 81.44 231.16 61.11 85.74 312.53 50.49 88.73 78.55
Gradnorm Huang et al. (2021) 134.85 76.20 68.51 305.28 75.61 75.57 405.68 65.57 77.46 78.55
SIRC Xia and Bouganis (2022) 59.57 63.38 86.70 259.93 77.99 80.50 320.35 56.93 87.45 78.55
LogitNorm Wei et al. (2022) 78.86 65.56 82.68 204.45 58.98 89.10 294.87 41.88 92.03 77.15
OpenMix Zhu et al. (2023a) 50.05 56.36 88.73 251.49 74.33 81.16 304.35 51.75 88.59 79.52
TAL 60.54 60.46 86.79 236.98 80.00 82.88 301.32 53.54 89.74 78.14
FMFP Zhu et al. (2022) 41.60 56.65 89.50 245.10 75.71 81.42 290.23 55.63 88.89 81.07
TAL w/ FMFP 46.71 61.16 87.49 209.80 61.18 88.32 270.90 40.63 92.50 80.98
DenseNet Huang et al. (2017) MSPHendrycks and Gimpel (2016) 79.06 63.62 85.75 257.69 77.86 79.77 333.86 57.58 87.64 74.62
Cosine Zhang and Xiang (2023) 82.36 63.65 84.72 250.14 77.01 81.48 332.18 53.82 88.28 74.62
Energy Liu et al. (2020) 108.80 72.79 78.73 240.11 75.65 82.81 341.87 58.95 86.94 74.62
MaxLogitHendrycks and Gimpel (2016) 105.98 70.20 79.60 240.06 75.65 82.86 339.68 56.69 87.39 74.62
Entropy Tian et al. (2022a) 79.94 65.41 85.46 250.68 75.54 81.34 330.88 53.97 88.30 74.62
Mahalanobis De Maesschalck et al. (2000) 159.95 95.30 64.54 208.53 54.83 89.23 358.90 63.29 84.93 74.62
Gradnorm Huang et al. (2021) 249.85 94.73 49.88 310.37 80.45 73.68 483.52 83.21 68.79 74.62
SIRC Xia and Bouganis (2022) 80.21 65.41 85.37 249.73 74.48 81.58 330.63 53.32 88.36 74.62
LogitNorm Wei et al. (2022) 109.80 73.28 78.90 216.63 60.21 87.27 330.07 48.43 89.50 73.80
OpenMix Zhu et al. (2023a) 64.11 59.52 87.29 267.67 79.70 78.27 328.27 58.83 86.79 77.03
TAL 81.73 63.87 84.39 219.48 67.16 86.43 309.96 45.16 91.12 75.14
FMFP Zhu et al. (2022) 56.67 62.45 87.72 258.04 77.11 79.31 314.57 59.81 87.37 77.96
TAL w/ FMFP 63.36 67.65 85.42 201.70 51.51 90.23 284.61 37.53 92.94 77.85
@@
ResNet50 MSP Hendrycks and Gimpel (2016) 72.73 63.95 86.18 301.27 46.01 87.21 351.26 49.64 86.99 76.13
Cosine Zhang and Xiang (2023) 102.98 69.93 79.49 298.35 50.64 87.54 359.74 54.43 86.17 76.13
Energy Liu et al. (2020) 118.66 76.33 75.81 279.16 35.64 90.47 351.93 43.69 87.74 76.13
MaxLogit Hendrycks and Gimpel (2016) 113.35 72.11 77.29 278.52 34.1 90.57 349.3 41.59 88.1 76.13
Entropy Tian et al. (2022a) 74.61 67.07 85.48 292.54 38.3 88.92 344.73 43.95 88.27 76.13
Mahalanobis De Maesschalck et al. (2000) 208.22 96.19 54.23 288.17 57.61 86.51 397.18 65.34 80.22 76.13
Residual Xia and Bouganis (2022) 238.18 97.01 49.0 316.1 57.77 83.89 431.12 65.55 77.12 76.13
Gradnorm Huang et al. (2021) 206.99 89.66 57.88 272.83 30.21 91.55 385.97 42.45 84.89 76.13
SIRC Xia and Bouganis (2022) 72.91 63.67 86.11 295.13 38.88 88.53 346.42 43.82 88.03 76.13
@@
ResNet50 MSP Hendrycks and Gimpel (2016) 72.73 63.95 86.18 272.93 59.27 87.72 326.85 60.19 87.42 76.13
Cosine Zhang and Xiang (2023) 102.98 69.93 79.49 255.91 68.67 89.85 326.97 68.92 87.81 76.13
Energy Liu et al. (2020) 118.66 76.33 75.81 235.88 37.67 94.22 318.89 45.27 90.60 76.13
MaxLogit Hendrycks and Gimpel (2016) 113.35 72.11 77.29 237.28 38.93 93.97 317.46 45.46 90.69 76.13
Entropy Tian et al. (2022a) 74.61 67.07 85.48 259.01 51.20 90.44 316.26 54.32 89.47 76.13
Mahalanobis De Maesschalck et al. (2000) 208.22 96.19 54.23 264.11 77.17 88.20 382.46 80.91 81.51 76.13
Residual Xia and Bouganis (2022) 238.18 97.01 49.00 282.46 81.30 85.09 409.89 84.39 77.99 76.13
Gradnorm Huang et al. (2021) 206.99 89.66 57.88 237.31 25.37 94.84 363.03 38.02 87.56 76.13
SIRC Xia and Bouganis (2022) 72.91 63.67 86.11 267.33 52.17 89.03 322.41 54.43 88.46 76.13

### v2

@@ 4.1 Comparisions with the State-of-the-art on CIFAR
ResNet110 MSP 99.83 67.49 84.07 293.44 83.41 74.55 376.42 66.92 84.00 72.01
Cosine 96.53 65.15 84.42 271.13 78.30 79.31 361.87 56.23 86.93 72.01
Energy 135.85 74.66 77.20 275.39 83.18 77.78 387.44 66.96 83.21 72.01
MaxLogit 133.19 72.33 77.96 275.85 82.53 77.73 385.81 65.08 83.56 72.01
Entropy 100.05 66.28 84.12 287.62 81.20 75.93 373.49 61.33 84.73 72.01
Gradnorm 114.21 73.48 80.41 263.49 72.70 80.55 368.55 58.74 85.74 72.01
Gradnorm 369.86 98.82 35.30 490.21 98.17 49.26 679.48 98.69 42.76 72.01
SIRC 100.56 66.37 84.01 287.93 81.03 75.90 374.12 61.29 84.65 72.01
LogitNorm 125.59 72.87 79.71 235.50 73.23 83.35 356.88 55.80 87.80 70.34
OpenMix 85.66 63.82 85.25 342.16 87.03 69.27 406.80 70.37 80.25 73.68
TAL 90.60 64.84 85.36 259.64 76.37 80.28 347.72 54.39 87.89 72.45
FMFP 69.83 62.17 87.15 284.13 81.77 74.98 345.37 62.99 84.86 75.18
TAL w/ FMFP 73.16 64.82 85.51 245.62 78.61 81.59 320.73 55.22 88.48 75.59
WRNet MSP 57.60 60.13 87.47 264.67 80.04 79.41 321.39 57.26 87.23 78.55
Cosine 56.68 58.65 87.56 279.53 81.60 77.91 334.04 57.96 86.08 78.55
Energy 66.33 65.64 84.73 257.78 78.93 80.83 323.27 58.70 87.07 78.55
MaxLogit 64.85 62.70 85.36 258.57 79.35 80.67 322.42 56.89 87.29 78.55
Entropy 59.31 62.78 86.80 260.92 79.50 80.26 320.66 56.92 87.40 78.55
Gradnorm 76.43 72.99 81.44 231.16 61.11 85.74 312.53 50.49 88.73 78.55
Gradnorm 134.85 76.20 68.51 305.28 75.61 75.57 405.68 65.57 77.46 78.55
SIRC 59.57 63.38 86.70 259.93 77.99 80.50 320.35 56.93 87.45 78.55
LogitNorm 78.86 65.56 82.68 204.45 58.98 89.10 294.87 41.88 92.03 77.15
OpenMix 50.05 56.36 88.73 251.49 74.33 81.16 304.35 51.75 88.59 79.52
TAL 55.41 58.43 88.41 245.36 79.48 81.29 303.18 54.62 89.22 78.14
FMFP 41.60 56.65 89.50 245.10 75.71 81.42 290.23 55.63 88.89 81.07
TAL w/ FMFP 44.21 58.77 88.40 224.56 71.32 85.31 278.28 47.14 91.05 80.98
DenseNet MSP 79.06 63.62 85.75 257.69 77.86 79.77 333.86 57.58 87.64 74.62
Cosine 82.36 63.65 84.72 250.14 77.01 81.48 332.18 53.82 88.28 74.62
Energy 108.80 72.79 78.73 240.11 75.65 82.81 341.87 58.95 86.94 74.62
MaxLogit 105.98 70.20 79.60 240.06 75.65 82.86 339.68 56.69 87.39 74.62
Entropy 79.94 65.41 85.46 250.68 75.54 81.34 330.88 53.97 88.30 74.62
Gradnorm 159.95 95.30 64.54 208.53 54.83 89.23 358.90 63.29 84.93 74.62
Gradnorm 249.85 94.73 49.88 310.37 80.45 73.68 483.52 83.21 68.79 74.62
SIRC 80.21 65.41 85.37 249.73 74.48 81.58 330.63 53.32 88.36 74.62
LogitNorm 109.80 73.28 78.90 216.63 60.21 87.27 330.07 48.43 89.50 73.80
OpenMix 64.11 59.52 87.29 267.67 79.70 78.27 328.27 58.83 86.79 77.03
TAL 73.28 60.90 86.51 250.78 75.46 81.35 325.70 52.05 88.73 75.14
FMFP 56.67 62.45 87.72 258.04 77.11 79.31 314.57 59.81 87.37 77.96
TAL w/ FMFP 60.52 63.65 86.57 224.62 63.43 86.01 296.75 45.10 90.93 77.85
@@
ResNet50 MSP 72.73 63.95 86.18 301.27 46.01 87.21 351.26 49.64 86.99 76.13
Cosine 102.98 69.93 79.49 298.35 50.64 87.54 359.74 54.43 86.17 76.13
Energy 118.66 76.33 75.81 279.16 35.64 90.47 351.93 43.69 87.74 76.13
MaxLogit 113.35 72.11 77.29 278.52 34.1 90.57 349.3 41.59 88.1 76.13
Entropy 74.61 67.07 85.48 292.54 38.3 88.92 344.73 43.95 88.27 76.13
Gradnorm 208.22 96.19 54.23 288.17 57.61 86.51 397.18 65.34 80.22 76.13
Residual 238.18 97.01 49.0 316.1 57.77 83.89 431.12 65.55 77.12 76.13
Gradnorm 206.99 89.66 57.88 272.83 30.21 91.55 385.97 42.45 84.89 76.13
SIRC 72.91 63.67 86.11 295.13 38.88 88.53 346.42 43.82 88.03 76.13
@@
ResNet50 MSP 72.73 63.95 86.18 272.93 59.27 87.72 326.85 60.19 87.42 76.13
Cosine 102.98 69.93 79.49 255.91 68.67 89.85 326.97 68.92 87.81 76.13
Energy 118.66 76.33 75.81 235.88 37.67 94.22 318.89 45.27 90.60 76.13
MaxLogit 113.35 72.11 77.29 237.28 38.93 93.97 317.46 45.46 90.69 76.13
Entropy 74.61 67.07 85.48 259.01 51.20 90.44 316.26 54.32 89.47 76.13
Gradnorm 208.22 96.19 54.23 264.11 77.17 88.20 382.46 80.91 81.51 76.13
Residual 238.18 97.01 49.00 282.46 81.30 85.09 409.89 84.39 77.99 76.13
Gradnorm 206.99 89.66 57.88 237.31 25.37 94.84 363.03 38.02 87.56 76.13
SIRC 72.91 63.67 86.11 267.33 52.17 89.03 322.41 54.43 88.46 76.13